OpenStudy (anonymous):

Condense 1/2log5^16-3log5^x+4log5^y

OpenStudy (anonymous):

Three condensing rules- log a + log b = log (ab) log a - log b = log (a/b) b log a = log (a^b)

OpenStudy (anonymous):

So now you use one, and plug your number into either formula.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."